### v13

@@ 2.2 Pareto optimal transfer mechanism with asymmetric information
This is because the transfer $\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda)$ in (8) is the solution to the problem (1) subject to $\tau(e_{t})\geq(u^{\prime})^{-1}(\lambda^{-1})-\overline{e}_{\lambda}$ for $t\in\mathbb{N}\cup\{0\}$, where $\lambda=\lambda_{0}$. In other words, since the transfer to the government is constant for individuals whose income is above the threshold $\overline{e}_{\lambda}$, they are indifferent between reporting and misreporting their true income.
With respect to the promise-keeping constraint, it holds because the expectation of the left-hand side of the inequality in (8) is greater than or equal to the promised level of lifetime utility $\overline{v}_{1}(\lambda)$:
$\mathrm{E}\big{[}(1-\beta)u(e+\tau^{*}(\min(e,\overline{e}_{\lambda})|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime})\big{]}\geq\overline{v}_{1}(\lambda).$
@@
An inter-period transfer mechanism $\mathcal{M}$ is represented by a sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ which satisfies (7) and the corresponding transfers, $\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$.
A sufficient condition for the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (6) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal transfer $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$. Write $F(\lambda)=\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))$ for $\lambda>0$.
@@
The sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (6) if
$-\frac{u^{\prime\prime}((u^{\prime})^{-1}(\lambda^{-1}))}{u^{\prime}((u^{\prime})^{-1}(\lambda^{-1}))}<\frac{\beta}{(1-\beta)(e^{M}-\mathrm{E}[e_{t}])}\quad\text{for}\quad\lambda>0.$ (9)
Proof.
@@
It suffices to show the proposition that $F(\lambda)<S$ for all $\lambda>0$.
@@
If condition (9) does not hold, there may be a case where an utility control variable $\lambda_{t}$ in $\mathcal{M}$ is defined in (7) instead of in (6), e.g. if the individuals’ utilities have constant absolute risk aversion, and if it is greater than $\beta/((1-\beta)(e^{M}-\mathrm{E}[e_{t}]))$. Since $F$ decreases in $\lambda$ and $\lim_{\lambda\to 0}F(\lambda)=\infty$, there must be a threshold $\overline{\lambda}$ below which $F(\lambda)\geq S$.
The inter-period transfer mechanism $\mathcal{M}$ is sequentially incentive compatible regardless of whether each $\lambda_{t}$ is defined in (6) or in (7). Furthermore, if $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (6), like the $\lambda$ mechanism of Marcet and Marimon [1992], $\mathcal{M}$ is sequentially efficient.
Proposition 1.
$\mathcal{M}$ is sequentially incentive compatible. Furthermore, if condition (9) holds, then $\mathcal{M}$ is sequentially efficient.
Proof.
@@
=(6)${=}\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(e|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}+v_{2}(\lambda_{t},e)-\overline{v}_{2}(\lambda_{t})$ )
@@
=(7)${=}\lambda_{t}\big{[}(1-\beta)u(e+\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t}))+\beta\overline{v}_{1}(\lambda_{t})\big{]}$
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{e,\overline{e}_{\lambda_{t}}\})-\mathrm{E}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
@@
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{\tilde{e},\overline{e}_{\lambda_{t}}\})-\mathrm{E}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
@@
The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
@@
We proceed to show that if condition (9) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms. By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible. By Lemma 1 since condition (9) holds, the sequence of utility control variables $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (6). Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period. Therefore $\mathcal{M}$ is a Pareto optimal transfer mechanism.
@@
It remains to prove that under condition (9) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism. Suppose, contrary to our claim, that there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$. Let $(v_{1}^{*},v_{2}^{*})$ be the present value achieved by $\Gamma$. Set $\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*})$. Since condition (9) holds for any $\lambda>0$, this $\lambda_{0}$ satisfies $F(\lambda_{0})<S$ as in the proof of Lemma 1. We may now use $(\lambda_{0},e)$ as the initial condition for $\mathcal{M}$ under condition (9). Then, by construction, each individual has the same present value for both contracts. Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$. However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under condition (9). ∎
Proposition 1 shows sequential efficiency of the mechanism $\mathcal{M}$ only for the case where condition (9) holds. However, we could relax this restriction by removing the assumption that the utility function is bounded from above. This assumption provides a sufficient condition to ensure Proposition 2, but it is not a necessary condition.
3 The inter-period transfer mechanism for each individual
The government’s approach of using the mechanism $\mathcal{M}$ successfully avoids an immiserization outcome. That is, the promised lifetime utility of each individual in $\mathcal{M}$ converges almost everywhere to a random variable with a finite expectation. To see this, we reinterpret $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}}$ as a sequence of random variables on $(\Omega,\mathcal{F},P)$. This is done recursively as follows. First we interpret $\overline{v}_{1}(\lambda_{1})$ as a random variable $\overline{v}_{1}(\lambda_{1}(e_{0}(\cdot)|\lambda_{0}))$ and let $\mathcal{F}_{1}\subset\mathcal{F}$ be the smallest $\sigma$-algebra induced by $\overline{v}_{1}(\lambda_{1})$. That is, $\mathcal{F}_{1}=\{\overline{v}_{1}(\lambda_{1})^{-1}(B)\ |\ B\in\mathcal{B}(\mathbb{R})\}$, where $\overline{v}_{1}(\lambda_{1})^{-1}(B)=\{\omega\in\Omega\ |\ \overline{v}_{1}(\lambda_{1}(e_{0}(\omega)|\lambda_{0}))\in B\}$. We see that $\overline{v}_{1}(\lambda_{1})$ is $\mathcal{F}_{1}$-measurable, $\overline{v}_{1}(\lambda_{1}):(\Omega,\mathcal{F}_{1})\to(\mathbb{R},\mathcal{B}(\mathbb{R})).$ Second, we consider $\overline{v}_{1}(\lambda_{n})$ as a random variable $\overline{v}_{1}(\lambda_{n}(e_{n-1}(\cdot)|\lambda_{n-1}))$ and define $\mathcal{F}_{n}\subset\mathcal{F}$ as the smallest $\sigma$-algebra induced by the product of the elements in $\{\overline{v}_{1}(\lambda_{i})\}^{n}_{i=1}$ such that
$\mathcal{F}_{n}=\{(\overline{v}_{1}(\lambda_{1})\times\dots\times\overline{v}_{1}(\lambda_{n}))^{-1}(B)\ |\ B\in\mathcal{B}(\mathbb{R}^{n})\},$
where
$(\overline{v}_{1}(\lambda_{1})\times\dots\times$ $\overline{v}_{1}(\lambda_{n}))^{-1}(B)$
=$=\Big{\{}\omega\in\Omega\ \Big{|}\ \big{(}\overline{v}_{1}(\lambda_{1}(e_{0}(\omega)|\lambda_{0})),\dots,\overline{v}_{1}(\lambda_{n}(e_{n-1}(\omega)|\lambda_{n-1}))\big{)}\in B\Big{\}}.$
We see that $\overline{v}_{1}(\lambda_{n})$ is $\mathcal{F}_{n}$-measurable, $\overline{v}_{1}(\lambda_{n}):(\Omega,\mathcal{F}_{n})\to(\mathbb{R},\mathcal{B}(\mathbb{R})).$
Proposition 2.
The sequence of promised lifetime utilities of individuals $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ converges to an integrable random variable.

### v15

@@ 2.3 Private information and mechanism design
This is because the transfer $\tau^{*}(\min\{e,\overline{e}_{\lambda}\}|\lambda)$ in (7) is the solution to the problem (1) subject to $\tau(e_{t})\geq(u^{\prime})^{-1}(\lambda^{-1})-\overline{e}_{\lambda}$ for $t\in\mathbb{N}\cup\{0\}$, where $\lambda=\lambda_{0}$. In other words, since the transfer to the planner is constant for individuals whose income is above the threshold $\overline{e}_{\lambda}$, they are indifferent between reporting and misreporting their true income.
Concerning the promise-keeping constraint, it holds because the expectation of the left-hand side of the inequality in (7) is greater than or equal to the promised level of lifetime utility $\overline{v}_{1}(\lambda)$:
$\mathbb{E}\big{[}(1-\beta)u(e+\tau^{*}(\min(e,\overline{e}_{\lambda})|\lambda))+\beta\overline{v}_{1}(\lambda^{\prime})\big{]}\geq\overline{v}_{1}(\lambda).$
3 Risk-shifting mechanism
This section builds and analyzes the transfer rule $\mathcal{M}$. We first show that each individual’s promised utility forms a bounded martingale, and then derive conditions under which the planner’s inter-temporal budget is non-negative.
@@ 3 Risk-shifting mechanism
(Risk-shifting mechanism) A mechanism $\mathcal{M}$ is a sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ and transfers $\tau_{t}(e_{t})=\tau^{*}(\min\{e_{t},\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ such that $\lambda_{t+1}$ is given by (6). Within each period $t$, consumption is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda_{t}^{-1})$; schocks affect only $\lambda_{t+1}$.
A sufficient condition for the sequence of promise $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ to always satisfy the recursive condition (5) is, as we see in the following lemma, that the Arrow-Pratt measure of the absolute risk aversion of $u$ at the optimal consumption $(u^{\prime})^{-1}(\lambda^{-1})$ is sufficiently low for all $\lambda>0$.
Assumption 1.
(Moderate risk aversion). For all $\lambda>0$,
$-\frac{u^{\prime\prime}}{u^{\prime}}((u^{\prime})^{-1}(\lambda^{-1}))<\frac{\beta}{(1-\beta)(e^{M}-\mathbb{E}[e_{t}])}.$ (8)
Interpretation. Arrow-Pratt risk aversion at the contract consumption level is smaller than a scaled income spread; the condition is mild when $\beta$ is close to one. Under Assumption 1, feasibility does not bind in the promise-update map, and the recursive rule (5) applies period by period.
Remark 2.
We could remove condition Assumption 1 by removing the assumption of bounded utility; dropping it would not violate sequential efficiency but would complicate the martingale proof. Boundedness is used to ensure that $\{\overline{v}_{1}(\lambda_{t})\}$ converges.
@@
Under Assumption 1 the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (5).
Proof: see Appendix A.
If the inequality (8) does not hold, there may be a case where a promised-utility state $\lambda_{t}$ in $\mathcal{M}$ is defined in (6) instead of in (5), e.g. if the individuals’ utilities have constant absolute risk aversion, and if it is greater than $\beta/((1-\beta)(e^{M}-\mathbb{E}[e_{t}]))$.
The risk-shifting mechanism $\mathcal{M}$ is sequentially incentive compatible regardless of whether each $\lambda_{t}$ is defined in (5) or in (6). Furthermore, if $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ is always defined in (5), like the $\lambda$ mechanism of Marcet and Marimon [1992], $\mathcal{M}$ is sequentially efficient.
Theorem 1.
The mechanism $\mathcal{M}$ defined by $(\tau^{\ast},\{\lambda_{t}\})$ is sequentially incentive compatible. Under Assumption 1, it is sequentially efficient.
Proof: see Appendix A.
3.1 Individual level: bounded promised utilities
This subsection shows that the sequence of promised utilities forms a bounded martingale and therefore converges, so immiseration cannot occur.
Let $(\mathcal{F}_{t})$ be the natural filtration generated by $\{\lambda_{s}\}_{s\leq t}$. Because (6) makes $\lambda_{t+1}$ measurable w.r.t. the $\sigma$-algebra generated by $\lambda_{t}$ and $e_{t}$, the process $\overline{v}_{1}(\lambda_{t})$ is adapted.
Theorem 2.
The promised utility process $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ induced by $\mathcal{M}$ is a bounded martingale and converges to an integrable random variable a.s. (no immiseration).
Proof: see Appendix A.
Because promised utility is a martingale with bounded support, the mechanism is a fair intertemporal lottery: expected lifetime utility is preserved ex-ante, but its realisation drifts neither up nor down on average.
3.2 Planner’s budget and sustainability
All claims in this subsection are sufficient (not necessary) conditions based on a second-order expansion. We derive sufficient conditions under which the cross-sectional budget is non-negative each period.
The convergence of individuals’ lifetime utilities to an integrable random variable with finite expectation does not yet ensure the sustainability of the $\mathcal{M}$ mechanism. This is because the planner predicts a declining trend in its budget for each individual in the long run, as we see in the following lemma. Let $\overline{\tau}_{n}$ denote the average transfer to an $n$-period-old individual (positive means the planner pays). Then $\overline{\tau}_{n}$ is:
$\overline{\tau}_{n}$ $=\mathbb{E}[\tau^{\ast}(e_{n}|\lambda_{n})]=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathbb{E}[e_{n}]\quad\text{ for }n\in\mathbb{N}\cup\{0\}.$
Lemma 2.
The average transfer for each cohort $\{\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ is an increasing sequence.
Proof sketch (details in Appendix B). For $n=1$, since $\mathbb{E}[\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
$\overline{\tau}_{1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathbb{E}[e_{1}]$
$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathbb{E}[e_{0}]=\overline{\tau}_{0}.$
Let $g_{t}(\lambda_{0})$ be the expected budget balance in period $t$ for the planner for a given initial promise $\lambda_{0}$. Since each individual’s probability of being alive is $\alpha\in(0,1)$, and from period one the next generation of continuum of individuals on the $(1-\alpha)$ interval is born in each period, $g_{t}(\lambda_{0})$ is given by
$g_{t}(\lambda_{0})$ $=\sum_{k=1}^{t}\alpha^{k}(\overline{\tau}_{k-1}-\overline{\tau}_{k})-\overline{\tau}_{0}.$
Here, a positive $g_{t}$ denotes a surplus (the planner collects funds). For the $\mathcal{M}$ mechanism to be sustainable, the budget balance in each period must be greater than or equal to zero. As we see in the following proposition, under a second-order (Taylor) approximation, if incomes have a symmetric distribution, a sufficient condition for the budget balance in period $t$ to be greater than or equal to be zero is that (i) the absolute risk aversion of $u$ is decreasing in consumption, (ii) the absolute risk aversion of the absolute risk aversion of $u$ is less than the absolute risk aversion of $u^{-1}$ and (iii) $\lambda_{0}$ is set so that the inequality $(\overline{\tau}_{1}-\overline{\tau}_{0})\alpha/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied.
Theorem 3.
In a stationary overlapping-generations society with i.i.d. income risk, if
  • (i)
    incomes have a symmetric distribution,
  • (ii)
    $u^{\prime\prime}/u^{\prime}$ is a decreasing function,
  • (iii)
    $(-u^{\prime\prime}/u^{\prime})\circ u^{-1}$ is a concave function and
  • (iv)
    initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ is satisfied,
then planner’s period-by-period balance $g_{t}(\lambda_{0})\geq 0$ for $t\in\mathbb{N}$ under a second-order (Taylor) approximation.
Proof: Second-order approximation shows decreasing differences. See Appendix C.
4 Numerical examples
An intuitive sufficient condition for $\lambda_{0}$ to satisfy the initial condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ specified in Theorem 3 is given by the second-order polynomial approximation as follows.
Corollary 1.
The condition $\alpha(\overline{\tau}_{1}-\overline{\tau}_{0})/(1-\alpha)\leq-\overline{\tau}_{0}$ in Theorem 3 holds with the second-order polynomial approximation if the initial promise $\lambda_{0}$ satisfies
where $\lambda_{E}=1/(u^{\prime}(\mathbb{E}[e_{t}]))$ denotes the promise corresponding to mean income report.
Proof: see Appendix D.
Figure 1: Utility Function
{singlespace*}
Notes: This figure illustrates a utility function (vertical axis) against consumption (horizontal axis). The segment $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$ appears in Corollary 1. It is the linear approximation of the length between the expected income and the consumption guaranteed in the mechanism M with $\lambda_{0}$.
The left-hand side of the inequality (9) is a linear approximation of the difference between the expected income and the consumption level in the mechanism $\mathcal{M}$ with the promise $\lambda_{0}$. The inequality (9) means that this length is greater than the square root of the weighted income variance. Figure 1 shows which part of the graph of a utility function corresponds to $\lambda_{0}(\overline{v}_{1}(\lambda_{E})-\overline{v}_{1}(\lambda_{0}))$. The lifetime utility $\overline{v}_{1}(\lambda_{0})$ in the mechanism $\mathcal{M}$ is higher than that in the state of autarky if the risk premium of the state of autarky is higher than the term in the right-hand side of the inequality (9).
A higher discount factor and higher probability of being alive will relax the sufficient condition of $\lambda_{0}$ for making the mechanism $\mathcal{M}$ sustainable. If $(u^{-1})^{\prime\prime\prime}$ is positive, the condition (9) may be unnecessarily strict.
We illustrate the mechanism with a CRRA utility, a symmetric discrete income distribution, and the benchmark parameter $r=\alpha=0.93$.
  • Utility $u(c)=\frac{1}{1+\gamma}(c+1)^{1+\gamma}$, $\gamma=-3$
  • Income support $E=\{0,1,2,3,4,5,6,7,8,9\}:=E_{s}$
  • Probabilities $P(e)$ (mirror-symmetric, approx. normal) $P_{s}$
  • Discount factor and survival rate $r=\alpha=0.93$
  • Initial state $\lambda_{0}=50\in[\lambda_{CE},133]$
The values follow Corollary 1: the lower bound $\lambda_{CE}=1/u^{\prime}\bigl{(}E[u(e_{t})]\bigr{)}=49.608$ ensures higher lifetime utility than autarky across all cohorts, and the upper bound $133$ is set to satisfy (9).
Figure 2 shows a series of sample means of lifetime utilities of $1000$ individuals of the same generation and corresponding transfer balances for $100$ periods for the above case. The initial promise is set $\lambda_{0}=50$. Consistent with Theorem 2, the utility path stays bounded while transfers $\{-\overline{\tau}_{n}\}_{n=0}^{100}$ fall, echoing Lemma 2.
Figure 2: Sample mean of lifetime utilities and transfer balances
Figure 3 shows a series of sample means of transfer balances for a society of $100$ individuals for $1000$ periods. In this society, each individual will be alive in the next period with probability $\alpha=0.93$, and newly born individuals enter the society, holding the total population $100$. As in Figure 2, each individual’s utility is $u(c)=(c+1)^{1+\gamma}/(1+\gamma)$ with $\gamma=-3$, and the parameters are the initial promise $\lambda_{0}=50$ and the discount factor $r=0.93$. The income set $E$ is $E_{s}$. The random numbers are generated with the same symmetric distribution $P_{s}$.
The series of sample means of transfer balances $\{g_{t}(\lambda_{0})\}_{t=0}^{1000}$ remains non-negative, and supports Theorem 3.
Figure 3: Sample means of budget balance $\{g_{t}(\lambda_{0})\}$
Remark 3.
Classic hidden-income contracts [e.g. Green, 1987, Thomas and Worrall, 1990] spread shocks within each period: truth telling is rewarded by a higher spot transfer, so lifetime utilities diverge over time and immiseration follows. Under our mechanism, by contrast, the realized income $e_{t}$ only determines next period’s promise $\lambda_{t+1}$; consumption in period $t$ is fixed at $c_{t}=(u^{\prime})^{-1}(\lambda^{-1}_{t})$ with probability $1$. The shock is therefore shifted one step forward—a transfer of risk to the future. This intertemporal insurance does make the average transfer profile $\{-\overline{\tau}_{n}\}_{n\in\mathbb{N}\cup\{0\}}$ fall with age, threatening the planner’s long-run budget. Theorem 3 shows the problem can be neutralized by choosing an initial promise $\lambda_{0}$ that gives early cohorts negative net transfers, so later positive transfers are prefunded. Finally, note a key difference from Marcet and Marimon [1992]: in their growth model, investment returns are themselves the incentive, and the contract is bilateral and non-competitive, which can depress the manager’s utility below autarky. Our pure-exchange setting has no such external scale effect, so every cohort attains a higher lifetime utility than autarky.
5 Dynamic-programming verification
This section verifies that the transfer rule $\tau^{\ast}$ solves the infinite-horizon Bellman equation and that $V$ is the corresponding value function (under condition (8) of Lemma 1).
The Bellman equation (DP) is given by
$V(\lambda_{t},t)=\sup_{\tau\in\Lambda(\lambda_{t})}\{(1-\beta)u(e_{t}+\tau(e_{t}|\lambda_{t}))+\beta\mathbb{E}_{t}[V(\lambda_{t+1},t+1)]\},$ (10)
where $\Lambda(\lambda_{t})$ is a set of measurable controls satisfying incentive and promise-keeping, and $\lambda_{t+1}$ is given by the function describing the change of the state variable $f:\mathbb{R}_{++}\times\mathbb{R}\times\mathbb{N}\cup\{0\}\to\mathbb{R}_{++}$ such that $f(\lambda_{t},\tau(\cdot|\lambda_{t}),t)=\lambda_{t+1}$ defined in (6).
Our candidate solution is $V$ in (11) with control $\tau^{\ast}$.
$V(\lambda_{t},t)=\mathbb{E}_{t}\big{[}u(e_{t}+\tau^{*}(e_{t}|\lambda_{t}))\big{]},\quad\text{for }\,\lambda_{t}\in\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}.$ (11)
We observe the following proposition.
Proposition 1.
With Assumption 1, $V$ in (11) solves the infinite-horizon Bellman equation, and $\tau^{*}$ maximizes the lifetime value $J$ in (2).
Proof sketch.
[…]
(complete proof in Appendix E).
6 Conclusion
This paper shows that the grim ”immiseration“ prediction of Green [1987] and Thomas and Worrall [1990] is not inevitable. A simple risk shifting rule that postpones risk can deliver near full within-period insurance while keeping promised utilities bounded. Provided the initial promise satisfies a transparent inequality, inter-generational cooperation sustains the mechanism, and every cohort attains higher lifetime utility than under autarky, with no within-cohort inequality.
Appendix A Proof of propositions 1 and 2
Proof of Lemma 1.
@@ Appendix A Proof of propositions 1 and 2
Write $F(\lambda)=\overline{v}_{1}(\lambda)+\lambda^{-1}\beta^{-1}(v_{2}(\lambda,e^{M})-\overline{v}_{2}(\lambda))$ for $\lambda>0$. It suffices to show the proposition that $F(\lambda)<S$ for all $\lambda>0$.
@@
Proof of Theorem 1.
@@
[…]
@@
[…]
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{e,\overline{e}_{\lambda_{t}}\})-\mathbb{E}_{\tau^{*}}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
@@
$\hskip 86.0pt+v_{2}(\lambda_{t},\min\{\tilde{e},\overline{e}_{\lambda_{t}}\})-\mathbb{E}_{\tau^{*}}[v_{2}(\lambda_{t},\min\{e_{t},\overline{e}_{\lambda_{t}}\})]$
@@
The last inequality follows from the fact that $\tau^{*}(\min\{e,\overline{e}_{\lambda_{t}}\}|\lambda_{t})$ corresponds to the solution of the problem (1) (planner’s static problem at given $\lambda$) subject to $\tau_{t}\geq(u^{\prime})^{-1}(\lambda^{-1}_{0})-\overline{e}_{\lambda_{0}}$ for $t\in\mathbb{N}$, where $\lambda_{0}=\lambda_{t}$.
@@
We proceed to show that if condition (8) holds, then $\mathcal{M}$ is Pareto optimal and not dominated by any other sequentially incentive-compatible mechanisms. By the first part of this proof, $\mathcal{M}$ is sequentially incentive compatible. By Lemma 1 since condition (8) holds, the sequence of promises $\{\lambda_{t}\}_{t\in\mathbb{N}\cup\{0\}}$ in $\mathcal{M}$ always defined in (5). Hence $\tau^{*}$ in $\mathcal{M}$ corresponds to transfers that solve the problem (1) in every period. Therefore $\mathcal{M}$ is a Pareto optimal transfer mechanism.
@@
It remains to prove that under condition (8) $\mathcal{M}$ is not Pareto dominated by any other sequentially incentive compatible mechanism. Suppose, contrary to our claim, that there exists a sequentially incentive compatible mechanism $\Gamma$ that Pareto dominates $\mathcal{M}$ for a given income state $e$. Let $(v_{1}^{*},v_{2}^{*})$ be the present value achieved by $\Gamma$. Set $\lambda_{0}=\overline{v}_{1}^{-1}(v_{1}^{*})$. Since condition (8) holds for any $\lambda>0$, this $\lambda_{0}$ satisfies $F(\lambda_{0})<S$ as in the proof of Lemma 1. We may now use $(\lambda_{0},e)$ as the initial condition for $\mathcal{M}$ under condition (8). Then, by construction, each individual has the same present value for both contracts. Since $\Gamma$ Pareto dominates $\mathcal{M}$, its Pareto dominance requires that $v_{2}^{*}>v_{2}(\lambda_{0},e)$. However, this contradicts the fact that solutions of $\mathcal{M}$ are Pareto optimal under the condition (8). ∎
Proof of Theorem 2.
$\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})(\omega)]=\overline{v}_{1}(\lambda_{n-1})(\omega)$
for given realized $\overline{v}_{1}(\lambda_{n-1})(\omega)$. We thus have the following equalities: For all $B\in\mathcal{B}(\mathbb{R})$ for all $n\in\mathbb{N}$
$\int_{B}\mathbb{E}\big{[}\overline{v}_{1}(\lambda_{n})\big{|}\overline{v}_{1}(\lambda_{n-1})(\omega)\big{]}dP_{\overline{v}_{1}(\lambda_{n-1})}$ $=\int_{B}\overline{v}_{1}(\lambda_{n-1})(\omega)dP_{\overline{v}_{1}(\lambda_{n-1})},$
and
$\int_{B}\mathbb{E}\big{[}\overline{v}_{1}(\lambda_{n})\big{|}\overline{v}_{1}(\lambda_{n-1})(\omega)\big{]}dP_{\overline{v}_{1}(\lambda_{n-1})}$ $=\int_{\{\overline{v}_{1}(\lambda_{n-1})(\omega)\in B\}}\overline{v}_{1}(\lambda_{n})dP$
$=\int_{\{\overline{v}_{1}(\lambda_{n-1})(\omega)\in B\}}\overline{v}_{1}(\lambda_{n-1})dP.$
From the last equality we have for $n\in\mathbb{N}$ $\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\mathcal{F}_{n-1}]=\overline{v}_{1}(\lambda_{n-1})$ a.e. $[P]$, and $\{\overline{v}_{1}(\lambda_{n})\}_{n\in\mathbb{N}\cup\{0\}}$ is a martingale as claimed.
Since $\{\overline{v}_{1}(\lambda_{n-1})\}_{n\in\mathbb{N}}$ is a martingale and $\sup_{n}\mathbb{E}[\overline{v}_{1}(\lambda_{n})^{+}]\leq\sup_{n}\overline{v}_{1}(\lambda_{n})^{+}<\infty$, from the submartingale convergence theorem, there is an integrable random variable $v^{\infty}$ such that $\overline{v}_{1}(\lambda_{n})\to v^{\infty}$ almost everywhere. ∎
Appendix B Proof of Lemma 2
Proof of Lemma 2.
We first consider the case $n=1$. For $n=1$, since $\mathbb{E}[\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0}))]=\overline{v}_{1}(\lambda_{0})$ and $u^{-1}$ is convex from Jensen’s inequality we have
$\overline{\tau}_{1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{1}(e_{0}|\lambda_{0})))]-\mathbb{E}[e_{1}]$
$\geq u^{-1}(\overline{v}_{1}(\lambda_{0}))-\mathbb{E}[e_{0}]=\overline{\tau}_{0}.$
For $n>1$, since $u^{-1}$ is a convex function, from Jensen’s inequality we see the result.
$\overline{\tau}_{n-1}$ $=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n-1}))]-\mathbb{E}[e_{n-1}]$
$=\mathbb{E}[u^{-1}(\mathbb{E}[\overline{v}_{1}(\lambda_{n})|\overline{v}_{1}(\lambda_{n-1})])]-\mathbb{E}[e_{n-1}]$
$\leq\mathbb{E}[\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))|\overline{v}_{1}(\lambda_{n-1})]]-\mathbb{E}[e_{n-1}]$
$=\mathbb{E}[u^{-1}(\overline{v}_{1}(\lambda_{n}))]-\mathbb{E}[e_{n}]$
$=\overline{\tau}_{n}\quad\text{ for }n=2,3,\cdots.$
Appendix C Proof of Theorem 3
The proof of Theorem 3 uses the following lemma.
Lemma C.1.
Under the assumptions (i), (ii), (iii) in Theorem 3, the second-order polynomial approximation of $\overline{\tau}_{n}-\overline{\tau}_{n-1}$ is a decreasing sequence.